College Graduation and Other State Characteristics (n = 50, k = 8)
omit
State ColGrad% Dropout EdSpend Metro% Age LPRFem Neast Seast West Midwest
AL 19.8 19.1 1221 89.2 37.4 55.8 0 1 0 0
AK 28.7 8.3 2187 74.7 33.9 65.6 0 0 1 0
AZ 27.9 14.3 1137 96.7 34.5 57.4 0 0 1 0
AR 17.5 18.6 1158 78.8 37.0 54.9 0 1 0 0
CA 30.4 19.7 1657 99.3 34.4 57.6 0 0 1 0
CO 35.4 10.8 1599 91.6 34.7 65.3 0 0 1 0
CT 36.9 10.0 2142 100.0 39.3 60.0 1 0 0 0
DE 25.6 13.2 1687 100.0 37.9 61.1 1 0 0 0
FL 25.5 13.2 1286 97.8 39.5 55.4 0 1 0 0
GA 26.9 14.3 1648 91.0 34.3 59.2 0 1 0 0
HI 30.6 12.7 1347 100.0 38.5 60.1 0 0 1 0
ID 25.9 10.9 1244 86.1 34.6 61.3 0 0 1 0
IL 29.5 12.7 2334 95.2 35.6 59.7 0 0 0 1
IN 22.5 12.8 1644 94.1 36.1 61.0 0 0 0 1
IA 24.5 10.1 1407 72.4 38.6 65.4 0 0 0 1
KS 30.4 8.6 1418 84.7 36.1 64.5 0 0 0 1
KY 19.0 21.1 1263 75.6 37.5 55.4 0 1 0 0
LA 19.7 19.8 1367 93.2 35.4 54.9 0 1 0 0
ME 24.2 12.8 1716 70.5 41.2 61.0 1 0 0 0
MD 36.3 13.0 1627 98.5 37.1 62.3 1 0 0 0
MA 36.8 12.5 1807 99.6 38.2 61.9 1 0 0 0
MI 24.6 11.5 1880 92.1 36.9 59.7 0 0 0 1
MN 34.3 7.3 1860 87.2 36.7 69.0 0 0 0 1
MS 21.9 20.1 1178 77.8 35.5 55.5 0 1 0 0
MO 24.9 14.6 1288 86.2 37.4 62.7 0 0 0 1
MT 25.4 7.9 1312 64.8 40.2 62.0 0 0 1 0
NE 25.1 10.3 1373 79.7 36.2 68.5 0 0 0 1
NV 23.5 13.4 1516 98.0 35.2 59.3 0 0 1 0
NH 32.8 8.1 1632 96.4 39.5 64.7 1 0 0 0
NJ 36.3 13.0 1903 100.0 38.0 58.4 1 0 0 0
NM 27.2 19.0 1608 95.9 36.2 57.5 0 0 1 0
NY 30.3 14.4 2046 97.7 37.5 56.2 1 0 0 0
NC 25.4 16.0 1233 91.7 36.2 58.8 0 1 0 0
ND 27.2 10.0 1316 69.2 39.1 67.6 0 0 0 1
OH 22.9 12.1 1667 95.5 37.6 60.4 0 0 0 1
OK 24.0 14.8 1163 84.2 36.5 57.6 0 0 0 1
OR 29.1 11.3 1384 96.1 37.0 59.0 0 0 1 0
PA 25.8 13.7 1617 96.9 39.7 58.1 1 0 0 0
RI 29.2 16.0 1594 100.0 38.4 61.7 1 0 0 0
SC 24.2 17.0 1422 93.5 37.1 59.5 0 1 0 0
SD 25.1 11.5 1329 71.3 37.0 69.4 0 0 0 1
TN 21.6 18.2 1079 89.5 37.3 57.4 0 1 0 0
TX 25.4 21.8 1696 93.8 33.2 58.2 0 0 1 0
UT 29.8 7.3 1319 94.8 28.5 62.7 0 0 1 0
VT 34.4 10.0 1948 73.7 40.7 65.8 1 0 0 0
VA 30.7 13.9 1631 89.1 37.2 60.8 0 1 0 0
WA 30.9 8.5 1561 96.4 36.7 61.2 0 0 1 0
WV 15.1 17.6 1538 75.0 40.7 49.1 0 1 0 0
WI 25.1 9.5 1792 85.9 37.9 66.6 0 0 0 1
WY 22.0 9.1 1896 71.5 39.1 65.3 0 0 1 0

Does the VIF on the LPR FEM coefficient indicate that mulitcollinearity is a problem?
Answer #1

The following problem is solved in minitab 2017.

Regression Analysis: ColGrad% versus Dropout, EdSpend, LPRFem, Metro%, Age, Neast, Seast, ...

The following terms cannot be estimated and were removed:
Midwest


Method

Categorical predictor coding (1, 0)


Analysis of Variance

Source DF Adj SS Adj MS F-Value P-Value
Regression 8 885.45 110.682 11.53 0.000
Dropout 1 18.76 18.757 1.95 0.170
EdSpend 1 18.70 18.704 1.95 0.170
LPRFem 1 81.48 81.480 8.49 0.006
Metro% 1 125.89 125.890 13.11 0.001
Age 1 2.74 2.738 0.29 0.596
Neast 1 108.81 108.807 11.33 0.002
Seast 1 12.91 12.915 1.34 0.253
West 1 37.11 37.115 3.87 0.056
Error 41 393.70 9.603
Total 49 1279.16


Model Summary

S R-sq R-sq(adj) R-sq(pred)
3.09879 69.22% 63.22% 55.22%


Coefficients

Term Coef SE Coef T-Value P-Value VIF
Constant -21.8 21.9 -0.99 0.327
Dropout -0.258 0.185 -1.40 0.170 2.64
EdSpend 0.00252 0.00181 1.40 0.170 1.45
LPRFem 0.548 0.188 2.91 0.006 3.17
Metro% 0.2036 0.0562 3.62 0.001 1.68
Age -0.146 0.273 -0.53 0.596 1.89
Neast
1 5.26 1.56 3.37 0.002 2.18
Seast
1 2.00 1.73 1.16 0.253 2.83
West
1 2.64 1.34 1.97 0.056 1.89


Regression Equation

Neast Seast West
0 0 0 ColGrad% = -21.8 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age

0 0 1 ColGrad% = -19.1 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age

0 1 0 ColGrad% = -19.8 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age

0 1 1 ColGrad% = -17.1 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age

1 0 0 ColGrad% = -16.5 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age

1 0 1 ColGrad% = -13.9 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age

1 1 0 ColGrad% = -14.5 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age

1 1 1 ColGrad% = -11.9 - 0.258 Dropout + 0.00252 EdSpend + 0.548 LPRFem
+ 0.2036 Metro% - 0.146 Age


Fits and Diagnostics for Unusual Observations

Obs ColGrad% Fit Resid Std Resid
8 25.60 32.64 -7.04 -2.41 R
44 29.80 31.80 -2.00 -0.97 X
50 22.00 27.93 -5.93 -2.21 R

R Large residual
X Unusual X

The VIF of LPRFem is 3.17.

Here as the VIF<5, then it can be concluded that multicollinearity is not a problem as such.